OpenStudy (alekos):

Here's a good one. Prove by induction that (n!)^1/n < (n+1)/2

OpenStudy (alekos):

Let's see if anyone can solve this gem!

OpenStudy (alekos):

\[\sqrt[n]{n!} \le \frac{ n+1 }{ 2 }\]

jhonyy9 (jhonyy9):

first step write for n=1 so than will get (1!)^(1/1) <= (1+1)/2 1! = 1 1^1 =1 1 <= 2/2 1 <= 1 so this is true now write for n=k and accept it true and on the next step write for k=k+1 and prove it true for n=k you get (k!)^(1/k) <= (k+1)/2 so this accept it true for k=k+1 will get ((k+1)!)^(1/(k+1)) <= ((k+1)+1)/2 ((k+1)!)^(1/(k+1)) <= (k+2)/2 so now you need prove this that is true for k=1 what will get ((1+1)!)^(1/(1+1)) <= (1+2)/2 (2!)^(1/2) <= 3/2 2! = 1*2 = 2 (2)^(1/2) = sqrt2 = 1,41 (1+2)/2 = 3/2 = 1,5 1,41 < 1,5 so what is true q.e.d. hope you understand it